### v2

# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

Abstract
Graphical model estimation from modern multi-omics data requires a balance between statistical estimation performance and computational scalability. We introduce a novel pseudolikelihood-based graphical model framework that reparameterizes the target precision matrix while preserving sparsity pattern and estimates it by minimizing an $\ell_{1}$-penalized empirical risk based on a new loss function. The proposed estimator maintains estimation and selection consistency in various metrics under high-dimensional assumptions. The associated optimization problem allows for a provably fast computation algorithm using a novel operator-splitting approach and communication-avoiding distributed matrix multiplication. A high-performance computing implementation of our framework was tested in simulated data with up to one million variables demonstrating complex dependency structures akin to biological networks. Leveraging this scalability, we estimated partial correlation network from a dual-omic liver cancer data set. The co-expression network estimated from the ultrahigh-dimensional data showed superior specificity in prioritizing key transcription factors and co-activators by excluding the impact of epigenomic regulation, demonstrating the value of computational scalability in multi-omic data analysis.
Keywords: Ultrahigh-dimensional molecular data, Multi-modal data, High-performance statistical computing, Graphical model selection, Pseudolikelihood, Communication-avoiding linear algebra
1 Introduction
Omics technologies are essential discovery tools for the analysis of complex biological samples such as tissues and biofluids. Individual omics modalities can generate high-dimensional data (large $p$) for characterizing complex biological samples in a high-throughput manner (large $n\ll p$), and recently for deconvoluting molecular profiles of isolated single cells or spatially connected micrometer-sized regions over tissue sections [37]. Given the unprecedented opportunity to assemble multi-omic data as fine-resolution descriptors of gene expression regulation, cellular metabolism and signal transduction in a biological system, clinical and basic research studies routinely employ a combination of omics platforms to characterize joint variation in the genome, epigenome, transcriptome, proteome and metabolome [12, 16].
In biomedical applications, the method of choice for multi-omic data interpretation is to map experimental data over well-characterized biological networks, where the data features are linked by a directed graph representing gene expression regulatory network, metabolic pathways or signaling cascades, or an undirected graph such as protein-6protein interaction network or gene co-expression network. Different networks facilitate human interpretation of molecular interactions in their respective biological context. However, a downside of this approach is that the network information biologically relevant to a given study may be unavailable in the literature, or the network information can be incomplete and thus fail to cover the molecules measured in a study. In other cases, the biological network of interest may be adaptive, and it is therefore susceptible to rewiring in specific conditions, rendering the available static network information inapplicable. For these reasons, it is often necessary to complement existing networks by estimating new ones directly from data so that a realistic snapshot of the relational structure among the variables is reflected in the downstream analysis or interpretation of the results.
In omics-scale analysis, biological network is typically represented as an adjacency matrix of data features based on marginal pairwise correlations. In the popular weighted gene co-expression network analysis [WGCNA, 41], for example, Pearson correlation-based network modules are captured first and the initial network is refined by topological analysis. Despite the success of WGCNA and other extensions in identifying gene communities from omics data, the adjacency of two molecular features as determined by marginal correlations can be a byproduct of shared regulatory factors rather than direct biological interaction. Hence it is natural to estimate the co-expression network underlying a gene expression dataset with a graph representing the conditional dependencies, e.g., partial correlations. The same logic extends to the inference of other types of biological networks from multi-omic data such as metabolic pathways and signaling cascades.
It is well known that the precision matrix, or inverse covariance matrix, of a multivariate distribution encodes partial correlations. The most popular high-dimensional precision matrix estimation approach is the graphical lasso [40, 11], which maximizes a multivariate Gaussian likelihood with an $\ell_{1}$-penalty. However, at the scale of modern multi-omics datasets, it is computationally infeasible to estimate desired networks using this strategy mainly due to the computational bottleneck, which is not easily resolved by using a more powerful computer. As we demonstrate later, our application data features close to 300,000 variables. On two high-performance computing (HPC) environments, the most scalable implementation of graphical lasso, BigQUIC [14], could complete the computation only for the trivial cases in which the computed precision matrix estimates were diagonal. Similarly, a fast implementation [30] of CLIME [4], another popular methodology, was also not applicable on a HPC machine with 192GB of memory when the dataset dimension exceeded 30,000. In the next section, we discuss in detail that the inherent nature of the optimization problems from these popular methods makes it difficult to scale up to modern omics-scale problems.
Feature screening procedures, which reduces the number of variables before the network estimation step, have been proposed as a remedy to address the lack of scalability of methods mentioned above [24, 42]. Although feature screening could be suitable when a moderate dimensionality reduction is necessary, an order of magnitude reduction (300,000 or more to 30,000 or fewer features) precludes functionally important molecules from being considered for network estimation for computational tractability rather than a biologically motivated rationale. It is also difficult to discern essential variables and dispensable ones a priori solely based on numerical criteria chosen by an analyst, hence there is a clear merit in estimating the model with all possible input variables included from the start.
Another category of graphical model structure learning approach is based on optimizing an $\ell_{1}$-penalized pseudolikelihood-based objective function [32, 18]. Although the structure of pseudolikelihood-based objectives can be advantageous from an optimization perspective [20], existing estimators do not possess the statistical properties desirable for multi-omics analyses, as we also discuss in the next section.
In summary, a statistical framework for partial correlation graph estimation scalable to contemporary multi-omics studies is currently lacking. To address this gap, we propose a new estimation framework called ACCORD that aims to strike a balance between computational scalability and statistical performance. In ACCORD, target precision matrix is reparameterized and estimated by minimizing the $\ell_{1}$-penalized empirical risk based on a new loss function. The associated optimization problem allows for a provably fast computation algorithm that is massively scalable, which is realized by a novel operator splitting algorithm and communication-avoiding distributed matrix multiplication. We then show that the ACCORD estimator possesses estimation consistency in $\ell_{1}$ and $\ell_{2}$ norms under standard high-dimensional assumptions and selection consistency under an irrepresentability condition. In this way, we balance between statistical performance and scalability of computation in massively large-scale settings.
We demonstrate that the HPC implementation of ACCORD, termed HP-ACCORD, scales well to handle dimensions up to one million. Leveraging on the scalability, we estimate a partial correlation network in the multi-omic liver cancer dataset from The Cancer Genome Atlas (TCGA). Using a combination of epigenomic and transcriptomic data sets with 285,358 variables in total, we successfully recapitulated local and global correlation structures of the variables within the same omics platform and identified a network of co-transcribed genes and DNA methylation events in upstream regulatory regions. More importantly, we show that the graph estimated from the ultrahigh-dimensional dual-omic data enabled us to identify bona fide transcription factors driving the co-expression network with greater specificity than and equivalent sensitivity to the alternative graph estimated from the transcriptomic data with 15,598 variables only. The analysis clearly highlights the merit of performing graph estimation in the whole feature space.
The organization of this paper is as follows. Section 2 explains the challenges that existing methods face when the scale of the problem becomes massively large. Section 3 introduces the ACCORD framework. In Section 4 we study the statistical properties of the ACCORD estimator. Section 5 is devoted to numerical experiments demonstrating the performance of the estimator and scalability of the algorithm. In Section 6, we showcase HP-ACCORD through graph estimation in a liver cancer data set with 285,358 multi-omic variables. The paper is concluded in Section 7.
2 Challenges
In this section we first describe the computational hurdles that current techniques for high-dimensional precision matrix estimation grapple with when dealing with extremely large scales.
Graphical lasso solves the convex optimization problem $\min_{\boldsymbol{\Theta}\in\mathbb{S}^{p}}\{-\log\det\boldsymbol{\Theta}+\operatorname{\bf tr}(\boldsymbol{S}\boldsymbol{\Theta})+\lambda\|\boldsymbol{\Theta}\|_{1}\}$, $\lambda>0$, in order to estimate the precision matrix $\boldsymbol{\Theta}^{*}$ of a $p$-variate zero-mean distribution. Here $\mathbb{S}^{p}$ is the space of $p\times p$ symmetric matrices, $\boldsymbol{S}=(1/n)\boldsymbol{X}^{T}\boldsymbol{X}$ is the sample covariance matrix, where $\boldsymbol{X}=[X_{1},X_{2},\cdots,X_{n}]^{T}$ with $X_{i}\in\mathbb{R}^{p}$ is the centered data matrix, and $\|\cdot\|_{1}$ is the vector $\ell_{1}$ norm. The estimate is fully characterized by the Karush-Kuhn-Tucker (KKT) optimality condition
$-\boldsymbol{\Theta}^{-1}+\boldsymbol{S}+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\|\boldsymbol{\Theta}\|_{1},$ (2.1)
where $\partial\|\boldsymbol{\Theta}\|_{1}$ denotes the subdifferential of the convex function $\boldsymbol{X}\mapsto\|\boldsymbol{X}\|_{1}$ at $\boldsymbol{\Theta}$. The vast literature on the algorithms for graphical lasso, e.g., [8, 11, 23, 13, 15], essentially reduces to how to solve the KKT equation (2.1) iteratively.
The computational culprit in (2.1) is the inverse of the $p\times p$ matrix variable $\boldsymbol{\Theta}$, which has to be computed every iteration. Inverting a $p\times p$ matrix using direct methods (e.g., Cholesky decomposition) costs $O(p^{3})$ arithmetic operations in general, which becomes prohibitive if $p$ is at the omics scale of a few hundreds of thousands. Storage requirements worsen the situation. The inverse $\boldsymbol{\Theta}^{-1}$ is in general not sparse, requiring $O(p^{2})$ memory space. For example if $p=300,000$, then roughly 700GB of memory is needed, which calls for distributed computation. Being inherently sequential, however, matrix inversion is difficult to parallelize or distribute, meaning that direct methods are not so much scalable even with high-performance computing (HPC) systems with distributed memory.
Consequently, most existing algorithms resort to employing an (inner) iterative method to compute $\boldsymbol{\Theta}^{-1}$. For instance, Friedman et al. [11] invoke a $p$-dimensional lasso regression solver $p$ times every (outer) iteration to compute one column of $\boldsymbol{\Theta}^{-1}$ at a time, which becomes prohibitively expensive in omics-scale problems. BigQUIC [14] updates $\boldsymbol{\Theta}$ by block coordinate descent and then computes $\boldsymbol{\Theta}^{-1}$ column by column, by storing only a small fraction of the $p$ columns of $\boldsymbol{\Theta}^{-1}$ in a cache and recomputing the missing columns on demand; the recomputation solves a fraction of the $p$ linear equations defining the inverse ($\boldsymbol{\Theta}\boldsymbol{X}=\boldsymbol{I}_{p}$) by using conjugate gradient. Since this strategy is only successful if the cache miss rate is low, BigQUIC partitions $\boldsymbol{\Theta}$ into a block matrix and permutes blocks to minimize the number of the off-block-diagonal elements. Therefore, the scalability of BigQUIC is limited by the maximum degree of the underlying graph. For more complex graph structures of interest, e.g., those that arise in multi-omics studies, computable dimensions are practically less than $100,000$; see §5.
CLIME minimizes $\|\boldsymbol{\Theta}\|_{1}$ subject to the constraint $\|\boldsymbol{S}\boldsymbol{\Theta}-\boldsymbol{I}_{p}\|_{\infty}\leq\lambda$ over $\mathbb{R}^{p\times p}$, where $\|\cdot\|_{\infty}$ is the vector $\ell_{\infty}$ norm. The optimization problem of CLIME can be decomposed into $p$ independent, $p$-dimensional linear programming (LP) problems, each solving a column of $\boldsymbol{\Theta}$. While LP is the most extensively studied convex optimization problem, commercial LP solvers like Gurobi or Cplex, which utilize interior-point or simplex methods, struggle with scalability when handling hundreds of thousands of variables [25]. Exacerbating the situation, for omics-scale CLIME we are required to solve hundreds of thousands of large-scale LP problems. This remains an almost unattainable task despite the independence. As a result, fastclime [30, 31], which could not handle more than 30,000 variables in our problem instance, remains one of the most scalable implementations publicly available.
As the most scalable pseudolikelihood-based method up to date, CONCORD [18] solves the $\ell_{1}$-penalized minimization problem
$\min_{\boldsymbol{\Theta}\in\mathbb{S}^{p}}\{-\log\det\boldsymbol{\Theta}_{D}+({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Theta}^{2}\boldsymbol{S})+\lambda\|\boldsymbol{\Theta}_{X}\|_{1}\},$ (2.2)
where $\boldsymbol{\Theta}_{D}$ denotes the diagonal and $\boldsymbol{\Theta}_{X}$ denotes the off-diagonal parts of $\boldsymbol{\Theta}$. It is a convex amendment of a non-convex pseudolikelihood-based method SPACE [32], which in turn extends the node-wise regression approach by Meinshausen and Bühlmann [27] to estimate $\boldsymbol{\Theta}^{*}$. [^2] CLIME can be understood as an extension of the latter in a different direction, with the aforementioned scalability bottleneck. The associated KKT condition is
$-\boldsymbol{\Theta}_{D}^{-1}+({1}/{2})(\boldsymbol{\Theta}\boldsymbol{S}+\boldsymbol{S}\boldsymbol{\Theta})+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\|\boldsymbol{\Theta}_{X}\|_{1}.$ (2.3)
Compared with that for graphical lasso (2.1), KKT condition (2.3) only involves the inverse of the diagonal matrix $\boldsymbol{\Theta}_{D}$, which is trivial to compute. While the second term in (2.3) costs two $p\times p$ matrix-matrix multiplications with $O(p^{3})$ operations, they are much easier to distribute and parallelize than matrix inversion. The sparsity of the optimization variable $\boldsymbol{\Theta}$ imposed by the $\ell_{1}$ penalty further reduces the complexity of the multiplication. Koanantakool et al. [20] leverage these features to achieve massive scalability in distributed memory HPC systems. This HP-CONCORD implementation can handle 320,000 dimensional data drawn from a random graph structure resembling multi-omic networks.
Despite its computational success, a major drawback of CONCORD is its lack of statistical consistency in estimating $\boldsymbol{\Theta}^{*}$, which is important in downstream tasks in multi-omic analyses as it carries information on the strength of network edges. The objective function in (2.2) defines a loss function $L(\boldsymbol{\Theta},x)=-\log\det\boldsymbol{\Theta}_{D}+\frac{1}{2}\operatorname{\bf tr}(x^{T}\boldsymbol{\Theta}^{2}x)$ for a sample $x\in\mathbb{R}^{p}$. The population risk $R(\boldsymbol{\Theta})=\operatorname{\bf E}L(\boldsymbol{\Theta},X)$ is minimized by a $\boldsymbol{\Theta}$ satisfying the equation $\boldsymbol{\Theta}_{D}^{-1}-\frac{1}{2}(\boldsymbol{\Theta}\boldsymbol{\Sigma}^{*}+\boldsymbol{\Sigma}^{*}\boldsymbol{\Theta})=\mathbf{0}$, where $\boldsymbol{\Sigma}^{*}={\boldsymbol{\Theta}^{*}}^{-1}$ is the covariance matrix of random vector $X$. Unfortunately, the true precision matrix $\boldsymbol{\Theta}^{*}$ does not minimize the risk unless $\boldsymbol{\Theta}_{D}$ equals the identity. This is a very restrictive assumption indeed, and likely does not hold in many applications. Consequently, the currently available result on the consistency of the CONCORD estimator [Theorem 2, 18] requires an accurate estimator $\widehat{\boldsymbol{\Theta}_{D}}$ of $\boldsymbol{\Theta}^{*}_{D}$ with a rate $\max_{i=1,\dotsc,p}|\widehat{\boldsymbol{\Theta}_{D}}_{,ii}-\boldsymbol{\Theta}^{*}_{D,ii}|=O_{P}(\sqrt{\log n/n})$. Similarly, SPACE also requires a separate estimator of $\boldsymbol{\Theta}_{D}^{*}$ with the same rate. Such a separate estimator is difficult to find in practice.
There is also an issue with the optimization method employed by HP-CONCORD. This CONCORD-ISTA algorithm [29] possesses an $O(1/t)$ rate of convergence, where $t$ is the number of iterations, and the convergence is measured in terms of the objective function value. In addition to the slow, sublinear convergence of the objective function, the variable iterate $\{\boldsymbol{\Theta}^{(t)}\}$ may converge to the minimizer $\hat{\boldsymbol{\Theta}}$ of (2.2) in an arbitrarily slow rate [2]. Since the algorithm has to be terminated within a finite number of iterations, statistical error of $\boldsymbol{\Theta}^{(t)}$ (say $\|\boldsymbol{\Theta}^{(t)}-\boldsymbol{\Theta}^{*}\|$) may remain quite large compared to that of $\hat{\boldsymbol{\Theta}}$ (say $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|)$ even after a large number of iterations.
3 The ACCORD Estimator and HP-ACCORD
3.1 ACCORD loss and estimator
For a matrix $\boldsymbol{\Omega}\in\mathbb{R}^{p\times p}$, consider the following loss function
$L(\boldsymbol{\Omega},x)=-\log\det\boldsymbol{\Omega}_{D}+(1/2)\operatorname{\bf tr}(x^{T}\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}x).$ (3.1)
The associated risk $R(\boldsymbol{\Omega})=\operatorname{\bf E}L(\boldsymbol{\Omega},X)=-\log\det\boldsymbol{\Omega}_{D}+\frac{1}{2}\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{\Sigma}^{*})$ is minimized by $\boldsymbol{\Omega}^{*}={\boldsymbol{\Theta}^{*}_{D}}^{-1/2}\boldsymbol{\Theta}^{*}$ since it is the unique solution to the equation $\nabla R(\boldsymbol{\Omega})=\mathbf{0}$ and $R$ is convex. The essential domain of $R$ on which $R$ is finite is $\operatorname{\bf dom}R=\{(m_{ij})\in\mathbb{R}^{p\times p}:m_{ii}>0,~{}i=1,\dotsc,p\}\cap\mathbb{R}^{p\times p}$. Then transformation $T:\boldsymbol{\Theta}\mapsto\boldsymbol{\Theta}_{D}^{-1/2}\boldsymbol{\Theta}$ from $\operatorname{\bf dom}R$ to itself is bijective with inverse $T^{-1}:\boldsymbol{\Omega}\mapsto\boldsymbol{\Omega}_{D}\boldsymbol{\Omega}$. Obviously $\boldsymbol{\Omega}^{*}=T(\boldsymbol{\Theta}^{*})$. In other words, the population risk is uniquely minimized by a one-to-one transformation $T$ of the true precision matrix $\boldsymbol{\Theta}^{*}$. If $\hat{\boldsymbol{\Omega}}$ minimizes the sample average of (3.1), then it consistently estimates $\boldsymbol{\Omega}^{*}$. Hence $T^{-1}(\hat{\boldsymbol{\Omega}})$ consistently estimates $\boldsymbol{\Theta}^{*}=T^{-1}(\boldsymbol{\Omega}^{*})$, as long as the dimension $p$ is fixed.
In a high-dimensional setting, we can expect minimizing an $\ell_{1}$-penalized sample average of (3.1) estimates $\boldsymbol{\Omega}^{*}=T(\boldsymbol{\Theta}^{*})$ consistently under the usual sparsity assumption on $\boldsymbol{\Theta}^{*}$:
$\hat{\boldsymbol{\Omega}}=\operatorname*{arg\,min}\nolimits_{\boldsymbol{\Omega}\in\mathbb{R}^{p\times p}}\left\{-\log\det\boldsymbol{\Omega}_{D}+(1/2)\text{tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})+\lambda\|\boldsymbol{\Omega}\|_{1}\right\};$ (3.2)
recall that $\boldsymbol{S}$ is the sample covariance matrix. An important point here is that the transformation pair $(T,T^{-1})$ preserves the sparsity pattern of their respective arguments: if $\omega_{ij}=[T(\boldsymbol{\Theta})]_{ij}$ with $\boldsymbol{\Theta}=(\theta_{ij})$, then $\omega_{ij}=\theta_{ij}/\sqrt{\theta_{ii}}$ and $\theta_{ij}=\omega_{ii}\omega_{ij}$. Therefore, $\omega_{ij}=0$ if and only if $\theta_{ij}=0$, and $\boldsymbol{\Omega}^{*}$ shares the sparsity pattern with $\boldsymbol{\Theta}^{*}$. The $\ell_{1}$ penalization on $\boldsymbol{\Omega}$ in (3.2) is thus justified. Since the downstream task is to analyze the partial correlations ($\rho^{*}_{ij})$, using the relation $\rho^{*}_{ij}=-\theta^{*}_{ij}/\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$ and $\theta^{*}_{ij}=\omega^{*}_{ii}\omega^{*}_{ij}$, we can consistently estimate $\rho^{*}_{ij}$ by the formula $\hat{\rho}_{ij}=-(1/2)(\hat{\omega}_{ij}/\hat{\omega}_{jj}+\hat{\omega}_{ji}/\hat{\omega}_{ii})$, in which $\hat{\boldsymbol{\Omega}}=(\hat{\omega}_{ij})$.
The loss (3.1) and estimator (3.2) resemble those of CONCORD in §2. The critical difference is that the argument $\boldsymbol{\Omega}$ is allowed to be asymmetric in our proposal. This simple change makes the estimator consistent (at least in the low-dimensional regime; high-dimensional results are presented in §4), a feature absent in CONCORD. For an obvious reason, we name the loss function (3.1) the ACCORD loss, and estimator (3.2) the ACCORD estimator, after Asymmetric ConCORD.
The KKT condition for the convex optimization problem for ACCORD (3.2) is
$-\boldsymbol{\Omega}_{D}^{-1}+\boldsymbol{\Omega}\boldsymbol{S}+\lambda\boldsymbol{Z}=\mathbf{0},\quad\boldsymbol{Z}\in\partial\left\|\boldsymbol{\Omega}\right\|_{1},$ (3.3)
which conserves the computational attraction of CONCORD; cf. (2.3). Furthermore, there is only one matrix multiplication in (3.3), suggesting simpler computation.
For the ACCORD estimator (3.2) to be well-defined, a solution to (3.3) needs to be unique. A matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ $(p\geq n)$ is said to have columns in general position if the affine span of any $n$ points $\{s_{1}X_{i_{1}},\cdots,s_{n}X_{i_{n}}\}$, for arbitrary signs $s_{1},\cdots,s_{n}\in\{-1,1\}$, does not contain any element of {±$\{\pm X_{i}:i\neq i_{1},\cdots,i_{n}\}$. General position occurs almost surely if $\boldsymbol{X}$ is drawn from a continuous distribution [35].
Theorem 3.1.
Suppose that the data matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ has columns in general position. Then, the ACCORD estimator (3.2) with $\boldsymbol{S}=(1/n)\boldsymbol{X}^{T}\boldsymbol{X}$ is unique.
3.2 ACCORD-FBS algorithm
For the computation of the ACCORD estimator, we use forward-backward splitting (FBS), also known as the proximal gradient method [6, 7]. Let us split the ACCORD objective function $f$ into $f=g+h$ where
$g(\boldsymbol{\Omega})=({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S}),\quad h(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+\lambda\|\boldsymbol{\Omega}\|_{1}.$ (3.4)
and apply a forward step for $g$: $\boldsymbol{\Omega}^{(t+1/2)}=\boldsymbol{\Omega}^{(t)}-\tau_{t}\nabla g(\boldsymbol{\Omega}^{(t)})$, and a backward (proximal) step for $h$: $\boldsymbol{\Omega}^{(t)}=\operatorname*{arg\,min}_{\boldsymbol{\Omega}\in\mathbb{R}^{p\times p}}\{h(\boldsymbol{\Omega})+\frac{1}{2\tau_{t}}\|\boldsymbol{\Omega}-\boldsymbol{\Omega}^{(t+1/2)}\|_{F}^{2}\}$ to obtain a sequence of optimization variables $\{\boldsymbol{\Omega}^{(t)}=(\omega_{ij}^{(t)})\}$. At the element level, we have the following closed-form iteration:
$\begin{split}\omega^{(t+1)}_{ii}&=(y-\tau_{t}\lambda+\sqrt{(y-\tau_{t}\lambda)^{2}+4\tau_{t}})/2,\quad y=\omega^{(t)}_{ii}-\tau_{t}[\boldsymbol{\Omega}^{(t)}\boldsymbol{S}]_{ii}\\ \omega^{(t+1)}_{ij}&=S_{\tau_{t}\lambda}(\omega^{(t)}_{ij}-\tau_{t}[\boldsymbol{\Omega}^{(t)}\boldsymbol{S}]_{ij}),\quad i\neq j,\end{split}$ (3.5)
where $S_{a}(x)=(|x|-a)_{+}\operatorname{sign}(x)$ is the soft-thresholding operator.
Theory of FBS ensures that with the choice of the step size $\tau_{t}\in(0,2/L)$, where $L=\sigma_{\max}(\boldsymbol{S})$ is the maximum singular value of $\boldsymbol{S}$, the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ converges to the target $\hat{\boldsymbol{\Omega}}$ [2]. The $L$ coincides with the global Lipschitz constant of $\nabla g(\boldsymbol{\Omega})=\boldsymbol{\Omega}\boldsymbol{S}$ in the Euclidean (Frobenius) norm. Backtracking line search adaptively finds a maximal step size $\tau_{t}$ for each iteration $t$ in such a way the convergence is guaranteed [3]. Algorithm 1, named ACCORD-FBS, describes both fixed step size and backtracking FBS algorithms for ACCORD. The step size satisfies $\tau_{t}\geq\min\{\tau_{0},\beta/L\}$ for $\beta\in(0,1)$ and the descent property $f(\boldsymbol{\Omega}^{(t+1)})\leq f(\boldsymbol{\Omega}^{(t)})$ is guaranteed by the majorization-maximization (MM) principle [22].
Algorithm 1 ACCORD-FBS
Input: sample covariance $\boldsymbol{S}\in\mathbb{R}^{p\times p}$, minimum step size $1/L$, initial step size $\tau_{0}$, line search parameter $10<\beta<$ 1, initial $\boldsymbol{\Omega}^{(0)}$
for $t$ in $0,1,2,\cdots$ do
    for $\tau_{t}$ in $\tau_{0},\beta\tau_{0},\beta^{2}\tau_{0},\cdots$ do
         $\nabla g(\boldsymbol{\Omega}^{(t)})\leftarrow\boldsymbol{\Omega}^{(t)}\boldsymbol{S}$triangleright Use Algorithm 2 for HP-ACCORD
         Update $\boldsymbol{\Omega}^{(t)}$ according to (3.5)
         $\boldsymbol{\Delta}\leftarrow\boldsymbol{\Omega}^{(t+1)}-\boldsymbol{\Omega}^{(t)}$
    end for if $g(\boldsymbol{\Omega}^{(t+1)})\leq g(\boldsymbol{\Omega}^{(t)})+\langle\boldsymbol{\Delta},g(\boldsymbol{\Omega}^{(t)})\rangle+\frac{1}{2\tau_{t}}\|\boldsymbol{\Delta}\|_{F}^{2}$ or $\tau_{t}\leq 1/L$
end for until converge
Output: estimate $\hat{\boldsymbol{\Omega}}\leftarrow\boldsymbol{\Omega}^{(t+1)}$
While the general convergence rate of FBS is $O(1/t)$ in the objective value unless the objective is strongly convex (which is not the case in ACCORD), we can nevertheless show that ACCORD-FBS converges linearly in both the objective and variable iterates.
Theorem 3.2.
Assume the condition for Theorem 3.1, and that the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ is generated by Algorithm 1 with the step size sequence $\{\tau_{t}\}$ satisfying either i) $\beta=1$, $\tau_{0}\in[\underline{\tau},\bar{\tau}]$, where $0<\underline{\tau}\leq\bar{\tau}<2/L$, or ii) $10<\beta<$ 1. Then, the objective value sequence $\{f(\boldsymbol{\Omega}^{(t)})\}$ converges to the minimum $f^{\star}=f(\hat{\boldsymbol{\Omega}})$ monotonically. Furthermore, if the initial iterate $\boldsymbol{\Omega}^{(0)}$ is chosen such that $f(\boldsymbol{\Omega}^{(0)})\leq\operatorname{\bf tr}(\boldsymbol{S})/2+\lambda p$, then the following holds
$\begin{split}f(\boldsymbol{\Omega}^{(t)})-f^{\star}&\leq\left(\frac{1}{1+2a\sigma}\right)^{t}[f(\boldsymbol{\Omega}^{(0)})-f^{\star}],\quad t\geq 0,\\ \|\boldsymbol{\Omega}^{(t)}-\hat{\boldsymbol{\Omega}}\|_{F}&\leq\left(\frac{1}{\sqrt{1+2a\sigma}}\right)^{t-1}\left(1+\frac{1}{a\sigma\sqrt{1+{1}/{(2a\sigma)}}}\right)\sqrt{{[f(\boldsymbol{\Omega}^{(0)})-f^{\star}]}/{a}},\quad t\geq 1,\end{split}$ (3.6)
where $a={1}/{\bar{\tau}}-{L}/{2}>0$, $\sigma=[4\kappa(1/\underline{\tau}+L)^{2}]^{-1}$ in case i, and $a=[2\tau_{0}]^{-1}$, $\sigma=[4\kappa(1/\tau_{\min}+L)^{2}]^{-1}$, $\tau_{\min}=\min\{\tau_{0},\beta/L\}$ in case ii. The constant $\kappa$ is explicit and depends only on $\boldsymbol{S}$.
Splitting (3.4) is crucial for establishing linear convergence in that the global Lipschitz constant $L<\infty$ of $\nabla g$ exists, and $01/L>$ 0 serves as the lower bound of the step size that guarantees the descent property. Following CONCORD-ISTA [29], we may also split $f$ into $\tilde{g}(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+({1}/{2})\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})$ and $\tilde{h}(\boldsymbol{\Omega})=\lambda\|\boldsymbol{\Omega}\|_{1}$ and call it ACCORD-ISTA. The problem with this splitting is that $\nabla\tilde{g}$ is not globally Lipschitz. The consequence is that the resulting step sizes satisfying the descent condition can be arbitrarily small, contributing to slow convergence of the iterates; see §2 and §5.
3.3 HP-ACCORD: HPC implementation of ACCORD
Algorithm 2 One-dimensionally distributed matrix multiplication (1DMM)
Input: Partition $\boldsymbol{\Omega}=[\boldsymbol{\Omega}_{1}^{T},\cdots,\boldsymbol{\Omega}_{P}^{T}]^{T}$, $\boldsymbol{S}=[\boldsymbol{S}_{1},\cdots,\boldsymbol{S}_{P}]$, where node $k$ holds $\boldsymbol{\Omega}_{k}$ and $\boldsymbol{S}_{k}$
for $k=1,\cdots,P$ simultaneously do
    for $j$ in $1\cdots,P$ do
         send $\boldsymbol{\Omega}_{k}$ to node $k-j$ and recv $\boldsymbol{\Omega}_{k+j}$ from node $k+j$
         Compute $\boldsymbol{\Omega}_{k+j}\boldsymbol{S}_{k}$
    end for
    $\boldsymbol{G}_{k}\leftarrow[(\boldsymbol{\Omega}_{1}\boldsymbol{S}_{k})^{T},\cdots,(\boldsymbol{\Omega}_{P}\boldsymbol{S}_{k})^{T}]^{T}$
end for
Output: $\boldsymbol{G}=[\boldsymbol{G}_{1},\cdots,\boldsymbol{G}_{P}]$, where node $k$ holds $\boldsymbol{G}_{k}$
For its simplicity and linear convergence rate, the ACCORD-FBS algorithm (Algorithm 1) has an advantage in scaling up to handle massive-scale data. Note that the main computational components of ACCORD-FBS are: 1) sparse-dense matrix multiplication in computing the gradient $\nabla g(\boldsymbol{\Omega}^{(t)})=\boldsymbol{\Omega}^{(t)}\boldsymbol{S}$ and 2) element-wise operations in computing (3.5); note that $\boldsymbol{\Omega}^{(t)}$ is sparse by construction. The latter is “embarassingly parallel.” The former can also be easily parallelized in shared memory systems such as those employing graphical processing units (GPUs). However, if the size of the data becomes massive so that it does not fit into the system memory, employment of a distributed memory system becomes necessary. In distributed computation, communication cost becomes a major factor of the performance.
Our distributed-memory HPC implementation of ACCORD-FBS, termed HP-ACCORD, iteratively conducts the sparse-dense matrix multiplication for the gradient step as a special case of the SpDM3 algorithm [20]. This algorithm, named one-dimensionally distributed matrix multiplication (1DMM), is summarized in Algorithm 2. In HP-ACCORD, $\boldsymbol{\Omega}$ and $\boldsymbol{S}$ are separated by columns (row-wise separation is also possible for $\boldsymbol{\Omega}$) and store each partition in different computational nodes. Then, the gradient $\nabla g(\boldsymbol{\Omega})$ is computed with 1DMM. While the standard method for distributed matrix-matrix multiplication on HPC systems is the scalable universal matrix multiplication algorithm [SUMMA, 36], this method partitions involved matrices by both rows and columns and allocates them across computational nodes, and completing each block of the product requires multiple rounds of broadcasting submatrices of both operands, which results in a major performance bottleneck. On the other hand, 1DMM sends and receives one-dimensional blocks of only one operand, i.e., $\boldsymbol{\Omega}$ in Algorithm 2. Because of this difference, the communication cost of Algorithm 2 is smaller than SUMMA since $\boldsymbol{\Omega}$ is sparse while $\boldsymbol{S}$ is dense [19]. Computation of the gradient can alternatively be conducted in two steps: $\boldsymbol{Y}=\boldsymbol{\Omega}\boldsymbol{X}^{T}$ and $\nabla g(\boldsymbol{\Omega})=(1/n)\boldsymbol{Y}\boldsymbol{X}$, each of which can be computed with 1DMM. This strategy is advantageous when $n$ is much smaller than $p$ [20].
3.4 Tuning
The choice of the regularization parameter $\lambda$ impacts the practical performance of the ACCORD estimator. At the omics scale, sample reuse methods such as cross validation or neighborhood selection [28] are ruled out since they incur a nontrivial number of expensive passes of computing the estimator. Following CONCORD and SPACE, which are also pseudolikelihood-based methods, we adopt a Bayesian information criterion (BIC)-type approach for tuning. Specifically, using the loss function part of (3.2), we choose $\lambda$ minimizing
$(2n)\left\{-\log\det\hat{\boldsymbol{\Omega}}_{D}+(1/2)\text{tr}(\hat{\boldsymbol{\Omega}}^{T}\hat{\boldsymbol{\Omega}}\boldsymbol{S})\right\}+\|\hat{\boldsymbol{\Omega}}\|_{0}\log n+4\gamma\|\hat{\boldsymbol{\Omega}}\|_{0}\log p,$ (3.7)
where $\|\hat{\boldsymbol{\Omega}}\|_{0}$ is the number of nonzero off-diagonal elements of the estimate $\hat{\boldsymbol{\Omega}}=\hat{\boldsymbol{\Omega}}(\lambda)$. The last term is taken from the extended BIC for graphical lasso [10], proposed to promote further sparsity; $\gamma\in(0,1]$ is a user-specified parameter. We may call quantity (3.7) an extended pseudo-BIC (epBIC).
3.5 Bias correction
Correcting the biases introduced by the $\ell_{1}$ penalization by refitting is a common practice [26, 5, 4]. Following these approaches, we propose the following second-stage refitting procedure for ACCORD. If we let the support of the ACCORD estimate $\hat{\boldsymbol{\Omega}}$ computed by solving (3.2) for an appropriate $\lambda$ (e.g., epBIC of (3.7)) be $\hat{S}_{\lambda}$, then we refit by computing
$\tilde{\boldsymbol{\Omega}}=\operatorname*{arg\,min}_{\boldsymbol{\Omega}:\;\boldsymbol{\Omega}_{\hat{S}_{\lambda}^{c}}=0}\left\{-\log\det\boldsymbol{\Omega}_{D}+(1/2)\text{tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})+\phi\lambda\|\boldsymbol{\Omega}\|_{1}\right\},$ (3.8)
where $0\leq\phi\leq 1$. The refitted estimator $\tilde{\boldsymbol{\Omega}}=(\tilde{\omega}_{ij})$ can be computed efficiently using Algorithm 1, by replacing the $\lambda$ with $\phi\lambda$ for $(i,j)\in\hat{S}_{\lambda}$ and with $\infty$ for $(i,j)\not\in\hat{S}_{\lambda}$.
4 Statistical Properties
In this section, we show that the ACCORD estimator can consistently estimate $\boldsymbol{\Omega}^{\ast}$, a one-to-one reparameterization of the true precision matrix $\boldsymbol{\Theta}^{*}$ in various measures, under appropriate conditions. All the results provided here are non-asymptotic. In addition to the vector $\ell_{1}$ and $\ell_{\infty}$ norms introduced in §1, we use $\vvvert\boldsymbol{M}\vvvert$ to denote an operator norm of matrix $\boldsymbol{M}=(m_{ij})$ induced by the underlying vector norm. In particular, $\vvvert\boldsymbol{M}\vvvert_{\infty}=\max_{\boldsymbol{x}\neq 0}\frac{\|\boldsymbol{M}\boldsymbol{x}\|_{\infty}}{\|\boldsymbol{x}\|_{\infty}}=\max_{i}\sum_{j=1}^{p}|m_{ij}|$. The Frobenius norm $\|\boldsymbol{M}\|_{F}$ of $\boldsymbol{M}$ is its vector $\ell_{2}$ norm. For an $l\times m$ matrix $\boldsymbol{M}$, $A\subset[l]:=\{1,\dotsc,l\}$, and $B\subset[m]$, we denote by $\boldsymbol{M}_{AB}$ the $|A|\times|B|$ submatrix of $\boldsymbol{M}$ taking the rows and columns of $\boldsymbol{M}$ with indices in $A$ and $B$, respectively.
4.1 Estimation error bounds
We first provide finite-sample estimation error bounds in vector $\ell_{1}$ and $\ell_{2}$ norms. Recall that a zero-mean random vector $Z$ is sub-Gaussian with parameter $\sigma$ if $\mathbb{E}[\exp(tZ)]\leq\exp({\sigma^{2}t^{2}}/{2})$ for all $t\in\mathbb{R}$. Let $\kappa_{\Omega^{\ast}}=\vvvert\boldsymbol{\Omega}^{\ast}\vvvert_{\infty}$.
Theorem 4.1.
Suppose the data matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ is composed of $n$ i.i.d. copies of zero-mean continuous random vector $X=(X_{1},\dotsc,X_{p})\in\mathbb{R}^{p}$ with covariance matrix $\boldsymbol{\Sigma}^{*}=(\Sigma^{*}_{ij})\allowbreak={\boldsymbol{\Theta}^{*}}^{-1}$ and each $X_{j}/\sqrt{\Sigma^{*}_{jj}}$ being sub-Gaussian with parameter $\sigma$. Also suppose that there exists $\alpha,\beta,\eta>0$ such that $\operatorname{\bf E}_{X}|\langle X,y\rangle|^{2}\geq\alpha$ and $\operatorname{\bf E}_{X}|\langle X,y\rangle|^{2+\eta}\leq\beta^{2+\eta}$ for any $y\in\mathbb{R}^{p}$ with $\|y\|_{2}=1$. If we let $S=\{(i,j)\in[p]\times[p]:\theta^{*}_{ij}\neq 0\}$ be the support of $\boldsymbol{\Theta}^{*}$ (hence of $\boldsymbol{\Omega}^{*})$, then there exist positive constants $\kappa,\allowbreak c_{0},\allowbreak c_{1},$ and $c_{2}$ that explicitly depends on $\alpha,\beta,\eta,\sigma,$ and $\max_{i\in[p]}\Sigma_{ii}^{*}$ such that $\lambda\!=64(1+4\sigma^{2})\kappa_{\Omega^{\ast}}(\max_{i\in[p]}\Sigma_{ii}^{*})\sqrt{n^{-1}\log p}$ in (3.2) yields
$\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{\ast}\|_{1}\leq 16\kappa^{-1}\lambda|S|\text{~{}and~{}~{}}\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{\ast}\|_{F}\leq 4\kappa^{-1}\lambda\sqrt{|S|}$
with a probability at least $1-c_{1}e^{-c_{2}n}-4p^{-2}$, provided that $n>\max\{(64c_{0}^{2}/\kappa^{2})|S|\log p,\allowbreak(1/16)\log p\}$.
Remark 4.1.
The consistency of $\hat{\boldsymbol{\Omega}}$ in terms of $\|\cdot\|_{1}$ and $\|\cdot\|_{F}$ is obtained if $\sigma^{2}$ and $\kappa_{\Omega^{\ast}}$ are bounded, at the rate of $O_{P}(\sqrt{n^{-1}|S|\log p})$. The latter is bounded if, e.g., the graph implied by $\boldsymbol{\Theta}^{*}$ has a finite maximum degree of $d=\max_{i\in[p]}|\{j\in[p]:\allowbreak\theta^{\ast}_{ij}\neq 0\}|$: since $|\theta^{*}_{ij}|\leq\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$, $\kappa_{\Omega^{\ast}}=\max_{i\in[p]}\sum_{j=1}^{p}|\theta^{*}_{ij}|/\sqrt{\theta^{*}_{ii}}\leq d\max_{i\in[p]}\sqrt{\theta^{*}_{ii}}\leq d/\sqrt{\lambda_{\min}(\boldsymbol{\Sigma}^{*})}$, where $\lambda_{\min}(\boldsymbol{\Sigma}^{*})\geq\alpha$ is the minimum eigenvalue of $\boldsymbol{\Sigma}^{*}$. This rate and the sample complexity of $n\gtrsim|S|\log p$ match those for the graphical lasso [39, Proposition 11.9].
4.2 Edge selection and sign consistency
With further assumptions, the elementwise ($\ell_{\infty}$) error bound of $\hat{\boldsymbol{\Omega}}$ can be controlled with which edge selection and sign consistency can be proved. The Hessian matrix of the population risk $R(\boldsymbol{\Omega})$ (see §3) with respect to the usual vectorization of $\boldsymbol{\Omega}$ at $\boldsymbol{\Omega}^{*}$ is
$\boldsymbol{\Gamma}^{\ast}=\boldsymbol{\Sigma}^{*}\otimes\boldsymbol{I}_{p}+(\boldsymbol{\Omega}_{D}^{\ast-1}\otimes\boldsymbol{\Omega}_{D}^{\ast-1})\boldsymbol{\Upsilon},$ (4.1)
where $\otimes$ is a Kronecker product and $\boldsymbol{\Upsilon}=\sum_{i=1}^{p}e_{i}e_{i}^{T}\otimes e_{i}e_{i}^{T}$. It is positive definite since so is $\boldsymbol{\Sigma}^{\ast}$. Let $S$ be the support of $\boldsymbol{\Theta}^{*}$ as stated in Theorem 4.1.
Assumption 4.1.
We further assume the following properties on the data distribution.
  1. (A1)
    The maximum degree $d=\max_{i\in[p]}\left|\{j\in[p]:\theta^{\ast}_{ij}\neq 0\}\right|=\max_{i\in[p]}\left|\{j\in[p]|\omega^{\ast}_{ij}\neq 0\}\right|$ of the graph implied by $\boldsymbol{\Theta}^{*}$ is bounded.
  2. (A2)
    (Irrepresentability condition) For some $\alpha\in[0,1)$, the matrix $\boldsymbol{\Gamma}^{*}$ satisfies
    $\vvvert\boldsymbol{\Gamma}^{\ast}_{S^{c}S}\boldsymbol{\Gamma}^{\ast-1}_{SS}\vvvert_{\infty}\leq 1-\alpha.$ (4.2)
Let us define the following associated quantities: $\kappa_{\Gamma^{\ast}}=\vvvert(\boldsymbol{\Gamma}^{\ast}_{SS})^{-1}\vvvert_{\infty}$, $\kappa_{\Sigma^{\ast}}=\vvvert\boldsymbol{\Sigma}^{\ast}\vvvert_{\infty}$, $\gamma_{1}=\vvvert\boldsymbol{\Omega}^{\ast-1}_{D}\vvvert_{\infty}$. Recall $\kappa_{\Omega^{\ast}}=\vvvert\boldsymbol{\Omega}^{\ast}\vvvert_{\infty}$.
Theorem 4.2.
Suppose the data matrix $\boldsymbol{X}\in\mathbb{R}^{n\times p}$ is composed of $n$ i.i.d. copies of zero-mean continuous random vector $X=(X_{1},\dotsc,X_{p})\in\mathbb{R}^{p}$ with covariance matrix $\boldsymbol{\Sigma}^{*}=(\Sigma^{*}_{ij})\allowbreak={\boldsymbol{\Theta}^{*}}^{-1}$ and each $X_{j}/\sqrt{\Sigma^{*}_{jj}}$ being sub-Gaussian with parameter $\sigma$. If further 4.1 holds, then for $\lambda=80\sqrt{2}(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\kappa_{\Omega^{\ast}}\alpha^{-1}\sqrt{(\tau\log p+\log 4)/n}$, $\tau>2$,
  1. (a)
    there holds $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{\ast}\|_{\infty}\leq 24\sqrt{2}(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\kappa_{\Gamma^{\ast}}\kappa_{\Omega^{\ast}}(1+10/\alpha)\sqrt{(\tau\log{p}+\log{4})/n}$;
  2. (b)
    the estimated support $\hat{S}=\{(i,j)\in[p]\times[p]:\hat{\omega}_{ij}\neq 0\}$ is contained in the true support $S$ and includes all edges $(i,j)$ with $|\omega^{\ast}_{ij}|>24\sqrt{2}(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\kappa_{\Gamma^{\ast}}\kappa_{\Omega^{\ast}}(1+10/\alpha)\sqrt{(\tau\log{p}+\log{4})/n}$,
with a probability no smaller than $1-p^{-(\tau-2)}$, provided $n>128(1+\sigma^{2})^{2}(\max_{i\in[p]}\Sigma_{ii}^{*})^{2}\delta^{-2}(\allowbreak\tau\log{p}+\log{4})$ where
$\delta=\min\left\{\frac{\min\left\{\frac{1}{3\gamma_{1}},\frac{1}{3\gamma_{1}^{3}\kappa_{\Gamma^{\ast}}},\frac{\kappa_{\Omega^{\ast}}}{3d}\right\}}{3\kappa_{\Gamma^{\ast}}\kappa_{\Omega^{*}}\left(1+{10}/{\alpha}\right)},\frac{2}{27\gamma_{1}^{3}\kappa_{\Gamma^{\ast}}^{2}\kappa_{\Omega^{\ast}}\left(1+{10}/{\alpha}\right)^{2}},8(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\right\}.$
Remark 4.2.
With other quantities held fixed, the sample size required to achieve the rate $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{*}\|_{\infty}=O_{P}(\sqrt{(\tau\log p)/n})$ is $n\gtrsim d^{2}\tau\log p$. It can be also shown that for data distributions with a bounded $4m$-th moment, the error rate is $O_{P}(\sqrt{p^{\tau/m}/n})$ for $n\gtrsim d^{2}p^{\tau/m}$; see Supplementary Material A.6. These sample complexities and rates of convergence match those of graphical lasso, obtained by Ravikumar et al. [33] under similar conditions.
Part (b) of Theorem 4.2 only states that the ACCORD estimator can exclude all false edges and find true edges with large enough $\omega_{ij}^{*}$’s. Exploiting it further, the sign consistency on all edges can be declared. For $\hat{S}=S$, let $\omega_{\min}=\min_{(i,j)\in S}|\omega_{ij}^{*}|$:
Theorem 4.3.
Assume the same conditions as Theorem 4.2. If the sample size satisfies
$n>128(1+4\sigma^{2})^{2}(\max_{i\in[p]}\Sigma_{ii}^{*})^{2}(\tau\log p+\log 4)/\min\{\omega_{\min}/[6\kappa_{\Gamma^{\ast}}\kappa_{\Omega^{\ast}}(1+10/\alpha)],\delta\}^{2}$
where $\sigma$ and $\delta$ are as defined in Theorem 4.2, then the event $\{\operatorname{sign}(\omega_{ij}^{*})=\operatorname{sign}(\hat{\omega}_{ij})$ for all $(i,j)\in S\}$ occurs with a probability no smaller than $1-p^{-(\tau-2)}$.
5 Numerical Experiments
5.1 Linear convergence of ACCORD-FBS
We provide empirical evidences of the guaranteed descent property and linear convergence exhibited by ACCORD-FBS (Algorithm 1). The merit of the novel operator splitting utilized by the latter is demonstrated by comparing it with ACCORD-ISTA (see the last paragraph of §3.2). Firstly, we generated two Erdos-Renyi graphs of size $p\!=\!1000$ with a sparsity level of $15\%$. To construct a ground truth precision matrix, we employed the following procedure. The edge weights were selected from a uniform distribution on $[0.5,1]$ and their signs were flipped with a probability of $0.5$. To ensure symmetry and positive definiteness, the resulting matrix was added to its transpose and its diagonal entries are set to be 1.5 times the absolute sum of the off-diagonal entries of the corresponding rows. This matrix was scaled by pre- and post-multiplying a diagonal matrix such that all the diagonal entries are equal to one. Finally, variation among the diagonal entries was introduced by performing another pre- and post-multiplications with a diagonal matrix with uniformly distributed entries on $[1,\sqrt{3}]$. Using this precision matrix, multivariate Gaussian data with a sample size of $n=500$ were generated.
ACCORD-FBS was investigated with two variants: one with a constant step size of $\tau={1}/{L}$ and the other employing backtracking line search with $\tau={1}/{L}$ as a lower bound. In contrast, a grid of constant step sizes, $\tau\in\{0.25,0.6,0.75\}$, along with backtracking, were chosen for ACCORD-ISTA to demonstrate its convergence behavior.
The convergence behavior of the iterate $\{\boldsymbol{\Omega}^{(t)}\}$ and the objective value $\{\ell(\boldsymbol{\Omega}^{(t)})\}$ is illustrated in Figure 1. Here, $\hat{\boldsymbol{\Omega}}$ denotes the final iterate obtained by executing ACCORD-FBS until termination, with the criterion of $\|\boldsymbol{\Omega}^{(t+1)}-\boldsymbol{\Omega}^{(t)}\|<10^{-15}$. In Figure 1, ACCORD-FBS exhibits linear convergence for both variants, while backtracking shows a faster rate. On the other hand, the convergence behavior of ACCORD-ISTA varies significantly across different constant step sizes, ranging from slow convergence to divergence. Moreover, the iterates from ACCORD-ISTA with backtracking encounters a plateau at an early stage. These observations underscore the challenges associated with the selection of an appropriate step size in ACCORD-ISTA.
Figure 1: Convergence comparison between ACCORD-FBS and ACCORD-ISTA.
5.2 Scalability of HP-ACCORD
We next investigate the scalability of HP-ACCORD (§3.3) using simulated data of dimension $p$ up to one million. At this scale, even simulating multivariate Gaussian samples becomes a nontrivial task [38]. We took a similar approach to the numerical experiments in Li et al. [23]: generate a sparse $p\times p$ unit lower triangular matrix $\boldsymbol{L}$ taking values in $[-1,1]$ and compute $\boldsymbol{y}=\boldsymbol{L}^{-T}\boldsymbol{x}$, $\boldsymbol{x}\sim N(\mathbf{0},\boldsymbol{I}_{p})$, by backsubstitution so that $\boldsymbol{y}$ follows $N(0,(\boldsymbol{L}\boldsymbol{L}^{T})^{-1})$. The location of off-diagonal non-zero entries of $\boldsymbol{L}$ were uniformly selected so that the graph implied by the precision matrix $\boldsymbol{L}\boldsymbol{L}^{T}$ had an average degree of $10.3$ for each $p$, and the maximum degree ranged from 42 to 66.
We employed the Nurion supercomputer at the Korea Institute of Science and Technology Information (KISTI) National Supercomputing Center (KSC) for the scalability experiment. Nurion is a Cray CS500 system with $8,305$ Intel Xeon Phi 7250 1.4GHz manycore (KNL) computational nodes with 96GB of memory per node, featuring 25.3 petaflops of peak performance. As a benchmark, we compared HP-ACCORD with BigQUIC. [^3] Available at https://bigdata.oden.utexas.edu/software/1035/. and fastclime. [^4] Available at https://cran.r-project.org/src/contrib/Archive/fastclime/. Note that fastclime is an R package with a single-core-oriented implementation of CLIME in C at its heart. Both BigQUIC and HP-ACCORD perform multi-core computations written in C++. However, BigQUIC’s scalability is limited to a single node as a shared-memory algorithm. In contrast, HP-ACCORD can run on multiple nodes simultaneously utilizing distributed memory in supercomputing environments (§3.3).
We report the results in Figure 2. Since the memory of a single KNL node (96GB) was not enough for the computation with dimensions greater than or equal to $100,000$, an appropriate number of nodes that can handle the input size data was used for HP-ACCORD, showcasing the scalability of the algorithm. The supercomputing center imposed a 48-hour restriction on the running time of a single job. Thus we only report accurate timing for the processes finished within this limit. In panel (a), the regularization parameter $\lambda$ was set large enough so that the resulting precision matrix estimate $\hat{\boldsymbol{\Theta}}$ becomes diagonal. In panel (b), $\lambda$ was adjusted so that the numbers of nonzero entries of $\hat{\boldsymbol{\Theta}}$ have a similar scale to their true precision matrices. Note that fastclime had to use the time near the limit to estimate a precision matrix to compute estimate for two $\lambda$’s for $p=10,000$, and it could not run with data that has $p=30,000$ and above due to the memory limit. Also, for the both panel, we could observe that HP-ACCORD outperforms BigQUIC in terms of computation time, even when only a single node is employed. By adopting multiple computational nodes, HP-ACCORD was able to deliver the estimate within the time budget in all scenarios. On the contrary, BigQUIC could not finish the process on time when the dimension exceeds 100,000 and the $\lambda$ is chosen so that the estimate is non-diagonal. Internally, BigQUIC tries to find a permuted partition $\{B_{1},\cdots,B_{k}\}$ of the $p$ coordinates using a graph clustering algorithm [17, 9] so that most of the coordinate updates are performed in the diagonal blocks $(B_{1},B_{1}),\cdots,(B_{k},B_{k})$. However, the size of each partition is limited due to the memory limit of a single computational node (about 20,000 in the Nurion environment in which each node has 96GB of memory), and clustering with such a limited block size bares many off-block diagonal edges in our simulated data. In consequence, the conjugate gradient method employed to compute the off-diagonal blocks are hardly skipped, causing the algorithm to run extremely slowly. Even when $\lambda$ is so large that the estimate should be diagonal, BigQUIC failed to complete the computation if the dimension reached one million.
(a) Large $\lambda$ (diagonal estimates)
(b) Moderate $\lambda$ (spase estimates)
Figure 2: Scalability comparison between HP-ACCORD and other methods. The horizontal red lines indicate the 48-hour limit of the supercomputer system.
5.3 Edge detection and estimation error
We examine the edge detection and parameter estimation performance of ACCORD on simulated data, comparing it with CONCORD, graphical lasso (GLASSO), SPACE, CLIME. Following the hub graph setting in Peng et al. [32], we generated a graph with 10 clusters each having 100 nodes and 90 edges and connected 100 randomly chosen node pairs from adjacent clusters. The clusters had either a hub network or scale-free graph structure. A hub network was generated by (1) constructing an Erdös-Rényi graph with 97 nodes and 45 edges, and (2) consecutively adding 3 hub nodes, each connected to random 15 nodes. A scale-free graph was generated so that the degree $k$ of the nodes has a probability $P(k)\sim k^{-2.3}$. We additionally considered a simple Erdös-Rényi graph with 1,000 nodes and 1,000 edges without the clustered structure. The partial correlation matrix corresponding to the graph was chosen so that the matrix has minimum eigenvalue of at least 0.2, and each nonzero entries have absolute value of 0.1 at least. Model selection was carried out using epBIC (3.7) for ACCORD and CONCORD for an evenly spaced grid of $\lambda$’s in logarithmic scale, while extended Gaussian BIC was used for GLASSO and CLIME. Although cross validation was used to select the $\lambda$ in the original paper of CLIME [4], we used BIC instead to avoid the repetitive parameter fitting. For SPACE, we used “BIC-type criterion” defined in Peng et al. [32]. For CONCORD, the matrix $\hat{\boldsymbol{\Theta}}$ minimizing (2.2) was treated as a precision matrix despite the possible inconsistency (see §2). This procedure was repeated 50 times. For ACCORD, we also considered debiasing procedure (3.8).
Graph Method AUPRC # TP edges # FP edges
Hub Network ACCORD 0.843 (0.011) 732 (18.6) 55 (17.0)
CONCORD 0.837 (0.011) 715 (18.1) 50 (16.5)
GLASSO 0.835 (0.010) 706 (20.5) 63 (11.3)
SPACE 0.813 (0.011) 768 (13.8) 321 (33.0)
CLIME 0.835 (0.010) 734 (13.8) 64 (17.5)
Scale-free ACCORD 0.882 (0.008) 810 (13.9) 77 (18.7)
CONCORD 0.882 (0.008) 797 (17.4) 67 (16.2)
GLASSO 0.874 (0.008) 808 (14.4) 110 (16.9)
SPACE 0.864 (0.008) 831 (13.8) 321 (33.0)
CLIME 0.879 (0.009) 813 (15.8) 85 (22.4)
Erdös-Rényi ACCORD 0.884 (0.009) 811 (14.5) 74 (16.9)
CONCORD 0.885 (0.008) 802 (14.7) 70 (14.5)
GLASSO 0.874 (0.009) 809 (16.0) 114 (12.7)
SPACE 0.868 (0.009) 836 (10.9) 369 (28.3)
CLIME 0.882 (0.009) 817 (13.5) 88 (19.3)
Table 1: Edge detection performance, mean (standard deviation) over 50 replications.
Table 1 reports the area under the precision-recall curve (AUPRC) and number of edges selected. In terms of AUPRC (and Matthews correlation coefficient; see Supplementary Material B), ACCORD performed slightly better than or similarly to other methods. The trend of the precision-recall curves did not particularly vary among replications. In all cases, the penalty selected from each selection method yielded reasonable number of selected edges. We could observe that BIC-type methods with the extended term drastically reduces false positive (FP) edges in the expense of few true positive (TP) edges.
Table 2 reports the total squared error of the estimated precision matrix $\boldsymbol{\Theta}$ and its reparameterization $\boldsymbol{\Omega}$ by ACCORD and CONCORD, along with their debiased refit. Compared to CONCORD, ACCORD estimates clearly showed better result for both $\hat{\boldsymbol{\Theta}}$ and $\hat{\boldsymbol{\Omega}}$ in terms of the estimate error. The debiased refit also clearly improves the estimating performance. For more details about the impact of the debiased refit, see Supplementary Material B.
Estimation Graph Total Squared Error ($\hat{\boldsymbol{\Theta}}$) Total Squared Error ($\hat{\boldsymbol{\Omega}}$)
ACCORD Biased Hub Network 194.6 (10.7) 73.7 (3.3)
Scale-free 202.6 (8.5) 68.1 (2.2)
Erdös-Rényi 201.0 (8.1) 69.7 (2.1)
Debiased Hub Network 37.7 (1.4) 19.7 (0.6)
Scale-free 36.2 (2.2) 16.5 (1.0)
Erdös-Rényi 35.4 (1.6) 16.3 (0.8)
CONCORD Biased Hub Network 392.3 (18.0) 108.0 (3.7)
Scale-free 433.6 (21.5) 107.9 (3.8)
Erdös-Rényi 437.5 (19.7) 110.8 (3.0)
Debiased Hub Network 206.6 (11.0) 44.2 (1.4)
Scale-free 226.6 (13.9) 43.6 (1.8)
Erdös-Rényi 231.0 (13.5) 44.8 (1.8)
Table 2: MSE of biased and debiased ACCORD and CONCORD in partial correlation, mean (standard deviation) over 50 replications.
6 Case Study: Integrative Analysis of Multi-Omics data in Liver Cancer
6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
In this section, we show that HP-ACCORD enables complex biological inference using ultrahigh-dimensional multi-omic data through the example of the LIver Hepatocellular Carcinoma (LIHC) cohort ($n$=365) of TCGA. In this analysis, we aim to identify the determinants of mRNA-level expression regulation of individual genes across the tumor samples and compare the strength of different regulatory drivers, particularly between the activity of transcription factors (TF) and the repressive potential of epigenomic regulators using partial correlations as the metric of independent contribution. Since the partial correlation between the mRNA expression level of a gene and the level of each regulatory factor given all other data features deconvolutes the strength of each source of direct regulation, this analysis, if computation is scalable to the entirety of the multi-omic data, will dissect relative contributions of individual regulatory mechanisms to the observed gene expression variation, and the causal graph mirroring the regulatory network is expected to be embedded within the reported (undirected) graph structure.
We used the LIHC dataset consisting of expression data of 15,598 protein-coding genes (mRNAs), expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands positioned in the upstream regulatory regions of genes. These are a subset of variables provided in the original data from the TCGA data portal [1], where only the DNA methylation probes in regulatory regions or gene bodies of annotated genes were included. Since DNA hypo- and hyper-methylation status tends to be correlated locally in genomic neighborhoods and those probesets for methylation status make up a majority of the variables in the data, the underlying graph is expected to reflect a sparse precision matrix with a large number of small block diagonals if the variables were ordered by chromosomes and by within-chromosomal genomic coordinates. Most causally implicated correlations between methylation probes and mRNAs are also likely to be local with the exception of DNA methyltransferase genes and the likes, and they would not alter the presumed network topology described above. The same logic also applies to mRNA-mRNA correlations, but these correlations can be local and global across genomic locations. However, these latter correlations tend to form block diagonal patterns by their upstream TF regulators and the target precision matrix will still remain within a structure amenable to consistent estimation.
We performed this analysis on two HPC systems: Nurion at KSC (see §5.2) and the Cori system at the National Energy Research Scientific Computing Center (NERSC), which is a Cray XC40 system with more than 2,000 computational nodes that are equipped with dual-socket 16-core Intel Xeon Processor E5-2698 (Haswell) and 128GB of memory for each node. On both systems, the shared-memory, multi-core implementation of BigQUIC was the only working alternative for comparison. However, BigQUIC was not able to complete the computation of the precision matrix for a wide range of the regularization parameter $\lambda$, except for those that yielded diagonal matrices; this observation is consistent with the results from the numerical studies in §5.2.
In the next two sections, we first describe the overall partial correlation patterns within and across the two omics data. Using the precision matrix estimate and partial correlations, we aim to tease apart the impact of active TF-driven co-regulation of target genes from that of DNA methylation-mediated gene expression repression in the promoter regions. To show the advantage of this analysis, we compare the mRNA-mRNA network from the two analyses, and show that the analysis accounting for the methylation-mediated regulation recovers experimentally validated regulatory actions of TFs with comparable sensitivity and better specificity. To the best of our knowledge, the effort to deconvolute the effects of two or more types of gene expression regulation using graphical models has not been attempted due to the computational bottleneck we address in this work.
6.2 Network structure of epigenomics and transcriptomics data
Figure 3: Summary of intra-modality and inter-modality correlations. (A) Barplots show the total number of edges with non-zero correlations in each of the six edge categories. (B) Estimated partial correlations against genomic distances between connected nodes. Gray dots represent all edges and they are shown in panels. Black dots show the data pertaining to the edge category in each panel. (C) Gene ontology terms enriched in the nodes (genes) with positive and negative methylation-mRNA partial correlations.
We estimated partial correlations by solving (3.2) at $\lambda=0.45$ selected by epBIC (3.7) and then applying the debiasing procedure (3.8). We extracted non-zero partial correlations to form edges and calculated the frequency of intra-modality and inter-modality pairs. Figure 3A shows the frequency of each type of inter- and intra-modality edges, with separate positive and negative edge counts.
Since the DNA methylation probes outnumbered both mRNA- and miRNA-level expression variables in the input data, the majority of the selected edges were the intra-modality correlations between methylation probes ($>$99%), which showed predominantly positive partial correlations (98%) even at long genomic distances. The second most frequent edges were between mRNAs and between miRNAs, where the vast majority showed positive partial correlations (96%), and the correlations were close to zero for the pairs associated with negative partial correlations. This observation clearly reaffirms that physically proximal protein coding genes and non-coding RNAs are often co-transcribed. Among the small number of non-zero inter-modality correlations, the most pronounced category was that of edges between DNA methylation probes and mRNA expression (total transcript level per gene); the signs of the partial correlations were evenly split between positive and negative (Figure 3A).
When we further examined the genomic distances between connected nodes across the six categories, positive inter-modality correlations were from the data features located on the same chromosomes and within 200 kilobase distance of one another (Figure 3B). The edges connecting mRNAs and miRNAs showed exclusively positive partial correlations within a 100 kilobase distance only, hinting at high local specificity of co-transcription. Last but not least, the edges connecting mRNAs and DNA methylation probes showed both positive and negative partial correlations. A closer examination revealed that most large negative partial correlations were between mRNAs and DNA methylation probes within a 1 kbp distance (from TSS), consistent with the established role of methylation for gene expression repression in regulatory regions.
When we further investigated biological functions enriched in the genes with negative correlation between DNA methylation and mRNA expression, those were mostly the genes encoding subunits of enzymes involved in small molecule metabolism and redox reaction regulation (Figure 3C), suggesting that epigenomic regulation actively co-represses the expression levels of metabolic enzymes in liver tumors.
6.3 Dissecting active regulatory contribution of TFs and repressive epigenomic regulation
Figure 4: (A) Comparison of the two partial correlation networks of mRNA expression data estimated with and without miRNA and DNA methylation data. (B) Comparison of partial correlation estimates for 110,787 edges appearing in the two mRNA-mRNA networks. (C) The network estimated with DNA methylation data. The able on the right side shows the biological processes and TFs enriched in each area.
Upon understanding the overall structure of the graph, we investigated the regulatory impact of DNA methylation on the expression levels of adjacent genes and dissected the regulatory contribution of TFs on their target gene expression and the repressive effects of hypermethylation in regulatory regions. To this end, we estimated another graph of mRNA expression variables using only the mRNA expression data as input, thereby establishing a reference network which does not exclude the impact of epigenomic regulation. We then compared the resulting mRNA-mRNA network against the one estimated earlier. For the comparison, we removed edges with absolute partial correlation smaller than 0.02 in the network visualization, an arbitrary threshold determined from the histogram of non-zero partial correlation estimates, to minimize the impact of false positive edges as mentioned in §5.
The new mRNA-mRNA network contained 104,599 edges at $\lambda=0.30$ based on epBIC. We observed that the graph estimated without the DNA methylation data, called unadjusted network hereafter, contained 84.5% of the edges from the graph estimated with the DNA methylation data, called adjusted network (Figure 4A). In those edges appearing in both networks, the magnitude of partial correlation estimates in the unadjusted network between mRNA pairs was greater than the one in the adjusted network, as expected (Figure 4B). Overall, the data clearly show that a significant portion of the conditional dependence relationship between two mRNA nodes can be explained away by DNA methylation levels, highlighting the contribution of epigenomic elements in the co-expression patterns.
Since the impact of methylation-driven co-repression of mRNA expression was accounted for, the adjusted network data allows us to infer active co-regulation driven by TFs and co-activators without potential confounding by epignomic regulation. We thus visualized the adjusted network using the Cytoscape software [34] and selected six subnetworks of high connectivity (Figure 4C). We performed hypergeometric probability-based tests for enrichment of biological functions and TFs potentially implicated in each subnetwork of genes. Indeed, the subnetworks showed highly specific enrichment of widely-known cancer associated biological processes, ranging from the extracellular matrix remodeling in the parenchymal environment, immune response and signal transduction, cell cycle and protein translation, and small molecular metabolism with previously well characterized TFs in various cellular processes.
By contrast, the unadjusted network showed a higher degree of connectivity between the nodes (visualization not shown due to lack of legibility) and the overall network lacked clear separation of subnetworks. Interestingly, the network had a large number of nodes with one or two edges connecting to others only. To investigate the quality of the two networks in terms of the recovery of well-validated TF-target relationships, we queried the selected mRNA-mRNA edges in the two networks against bona fide TF-target regulatory network compiled from credible sources [21]. Upon mapping the two mRNA-mRNA networks, 32% and 30% of the edges involved genes encoding a TF or a co-activator in the adjusted and unadjusted networks, respectively. However, while 19% of the reported edges were experimentally validated TF-target pairs in the adjusted network, only 9% of the edges were validated in the unadjusted network. Given the unadjusted network is about 2.6 times larger than the adjusted network, the result implies that the adjusted network recovers the known TF-target regulatory relationships with comparable sensitivity and greater specificity.
In summary, the assessment clearly shows that the network derived from the ultrahigh-dimensional dual-omic data represents more robust relational hypotheses than the network derived from the transcriptomic data alone. With the capability to estimate conditional dependence structure from ultrahigh-dimensional data, spurious associations can be screened out and the results enables subsequent inference of regulatory relationships, which would have been otherwise impossible.
7 Discussion
In discovery-oriented clinical and basic research applications, technological advances in molecular modalities have steadily increased the number of variables to analyze. Typical dimensionality of data sets exceeds tens or hundreds of thousands, as shown in the motivating dual-omic dataset. The latest technological advances also push sample sizes to thousands, best evidenced by the number of cells analyzed using single-cell resolution omics technologies. When making inference from these truly big data, striking a balance between statistical performance and computational scalability is therefore essential.
The analysis in §6 is a testament to this trade-off. Although the validity of individual edges should be confirmed by experimental means, we verified that the mRNA co-expression network estimated from the ultrahigh-dimensional dual-omic data was largely nested within the network obtained from mRNA expression variables alone, not deviating to a completely distinct result. Moreover, the former network excluding the epigenomic regulation allowed us to recover bona fide TF-target relationships validated in human cells with greater specificity and without loss of sensitivity.
The analysis of dual-omic data presented in this paper would have been impossible without the scalable computation for estimating the underlying graphical model. In an era with new omics modalities added to the multi-omic repertoire and declining cost to produce such data, graphical model estimation in ultrahigh dimensional space will be in increasing demand and computational scalability will thus be as essential as ever. We believe HP-ACCORD paves the way to meeting the demand for graphical model estimation in the modern multi-omic studies and encourages development of related statistical frameworks having scalability in their backbone.